# The Loser’s Curse and the Critical Role of the Utility Function

### v3

@@ Abstract
A longstanding question in the judgment and decision making literature is whether experts, even in high-stakes environments, exhibit the same cognitive biases observed in controlled experiments with inexperienced participants. In their seminal work, Massey and Thaler (2013) provide a notable example of bias and irrationality in expert decision making: general managers’ behavior in the National Football League draft pick trade market. They argue that general managers systematically overvalue top draft picks, which generate less surplus value on average than later first-round picks, a phenomenon known as the loser’s curse. Their conclusion hinges on the assumption that general managers should use expected surplus value as their utility function for evaluating draft picks. This assumption, however, is neither explicitly justified nor necessarily aligned with the strategic complexities of constructing a National Football League roster. In this paper, we challenge their framework by considering alternative utility functions, particularly those that emphasize the acquisition of transformational players––those capable of dramatically increasing a team’s chances of winning the Super Bowl. Under a decision rule that prioritizes the probability of acquiring elite players, which we construct from a novel Bayesian hierarchical Beta regression model, general managers’ draft trade behavior appears rational rather than systematically flawed. More broadly, our findings highlight the critical role of carefully specifying a utility function when evaluating the quality of decisions.
@@ 1 Introduction
A longstanding question in the judgment and decision making literature is whether experts, when making high-stakes decisions, exhibit the same cognitive biases and irrational behaviors documented in studies involving inexperienced participants in low-stakes environments. Critics of this literature argue that biases may be overstated, as most studies do not analyze decision making in real-world, high-stakes settings. Massey and Thaler, (2013) provide a notable exception, applying behavioral economic principles to a context––the National Football League (NFL) draft––where decision makers, namely NFL general managers, are highly experienced, financially incentivized, and have access to extensive data and learning opportunities.
Massey and Thaler’s study has had a significant impact not just in sports analytics but in the broader decision making literature, reinforcing the notion that systematic biases can persist even in expert settings. They argue that general managers systematically overpay for top draft picks despite evidence that these selections generate less “surplus value”––the difference between the value of a player’s performance and his cost––than later first-round picks, on average. This so-called “loser’s curse,” widely cited as evidence of irrational decision making, is the subject of their ultimate striking conclusion: “top draft picks are significantly overvalued in a manner that is inconsistent with rational expectations and efficient markets, and consistent with psychological research” (Massey and Thaler,, 2013).
@@ 1.1 Overview
Players entering the NFL are allocated to 32 teams through an annual draft. The draft consists of seven rounds, in which teams take turns selecting players in reverse order of the previous season’s win-loss records. The team with the worst record picks first, while the reigning Super Bowl champion selects last (with some minor exceptions). There are $256$ total draft picks because the NFL adds $32$ additional “compensatory free agent” draft picks––creating a hidden eigth round––that allow teams to replace players who they lost in free agency.111 https://operations.nfl.com/journey-to-the-nfl/the-nfl-draft/the-rules-of-the-draft Trading draft picks, involving the exchange of current and future draft selections as well as active players, is a fundamental component of constructing an NFL roster. Teams typically engage in these trades to secure specific players or improve positional depth. For example, in 2013 the Miami Dolphins traded the $3^{rd}$ overall pick to the Oakland Raiders in exchange for the $12^{th}$ and $42^{nd}$ overall picks.222 https://bleacherreport.com/articles/1618345-miami-dolphins-select-dion-jordan-with-no-3-pick-in-trade-with-oakland-raiders Trades like this raise fundamental questions about the valuation of draft picks. How did the Dolphins and Raiders determine that their trade was mutually beneficial? More broadly, if a team seeks to acquire a specific draft pick, what constitutes an appropriate offer? If presented with a trade proposal involving multiple draft picks, how should a team evaluate its value and decide whether to accept or reject the exchange? A rigorous analysis of draft pick valuation is essential to understanding the trade dynamics that shape NFL roster management.
@@
More recently, in their seminal work, Massey and Thaler, (2013) developed the first rigorous data-driven draft position value curves, which are visualized in Figure 1. The solid line in Figure 1a denotes expected performance value, which quantifies the expected on-field performance of players drafted at a given position, expressed on a salary-equivalent scale. The dot-dashed line reflects first contract compensation, which denotes the cost of drafted players during their first contract. A player’s first contract compensation is essentially deterministic, as it is dictated by draft position. [^4] Drafted rookies may negotiate certain contract elements, such as performance bonuses or injury guarantees. However, this compensation is often much lower than the value of the player’s on-field performance, leading to what is known as first contract surplus value. The dashed lines in Figures 1a and 1b reflect expected surplus value, which measures the expected difference between a player’s first contract performance and his cost. Finally, the solid line in Figure 1b is the trade market curve, a draft position value curve that best fits the dataset of all observed trades involving only draft picks. This curve reflects the draft value implied by actual trade market behavior.
@@
Massey and Thaler identified two key insights. First, they discovered the loser’s curse, visualized in Figure 1a. Teams with the worst win-loss records in the previous season—the losers—receive the top draft picks, as expected performance value declines monotonically and convexly throughout the draft. However, these teams are cursed because top draft picks generate less surplus value on average than mid-to-late first-round picks, which typically belong to better teams. Expected surplus value (i.e., expected performance above cost) increases throughout the early first round, peaking in the middle of the round. Second, Massey and Thaler demonstrated that, if one adopts the expected surplus value utility function, the draft pick trade market is highly inefficient, as shown in Figure 1b. The trade market curve is much steeper than the expected surplus value curve and differs in shape. Teams consistently overpay to trade up in the draft, investing excessive draft capital to select “top draft picks [that] are significantly overvalued” (Massey and Thaler,, 2013). In contrast, trading down and accumulating additional draft picks is an effective strategy to maximize surplus value. For example, consider the 2013 draft trade between the Dolphins and the Raiders. Using expected surplus value,555 This calculation is based on our expected surplus value curve, visualized in Figure 2 in Section 3, as the exact values from Massey and Thaler’s curve are unavailable. the Raiders’ draft capital increased by 135%. The Dolphins traded two picks collectively worth 2.40 points (1.51 and 0.89, respectively) in exchange for a single pick worth 1.02. Notably, the $12^{th}$ overall pick had a higher expected surplus value than the $3^{rd}$ overall pick (i.e., $1.51>1.02$). Although the Dolphins won the trade according to Jimmy Johnson’s draft value chart, the Raiders accumulated substantially more expected surplus value.
@@
Massey and Thaler’s findings, summarized in Figure 1, are corroborated by recent data666 That Massey and Thaler’s results replicate in recent data was originally noticed as part the development of a forthcoming paper Avery et al., ming (see Figure 2 in Section 3). Despite the widespread recognition of their work over an extended period, the draft strategies employed by general managers in the NFL have remained largely unchanged. This persistence raises an important question: why has there been no apparent adjustment in decision making? One possible explanation––examined in the forthcoming paper Avery et al., ming––is that NFL general managers are resistant to change and continue to engage in suboptimal trade practices due to cognitive biases or misjudgment. However, this perspective prompts further inquiry: are these executives, responsible for managing multi-billion-dollar franchises, fundamentally misguided, or is there a crucial element missing from existing analyses? This paper explores the latter hypothesis, proposing an alternative explanation. Specifically, we argue that there may be a sound economic rationale underlying observed trade behavior. It is plausible that general managers are employing a different, yet entirely rational, utility function that does not rely solely on the expected future surplus value of a player at a given draft position.
@@
Perhaps the overarching strategic objective of NFL teams is to win the Super Bowl––a goal consistently emphasized by team owners, general managers, coaches, and players. As Kansas City Chiefs quarterback Patrick Mahomes succinctly stated, “the only goal is to win the Super Bowl.” [^7] https://x.com/_MLFootball/status/1826003252460269907?t=JvStTWyM92jEe21LxqOTPw&s=19 Massey and Thaler’s conclusions are predicated on the assumption that general managers, in pursuit of this objective, should adopt expected surplus value as their guiding utility function. However, while their analysis relies on this assumption, they do not explicitly acknowledge the primacy of winning the Super Bowl or establish a direct link between maximizing expected surplus value and achieving that goal. While it is possible that maximizing expected surplus value contributes to winning, alternative utility functions may yield entirely different strategic conclusions. This paper develops that possibility, exploring the idea that Super Bowl success hinges on acquiring transformational players––those capable of profoundly altering a team’s trajectory––rather than optimizing expected surplus value. One example is the selection of Patrick Mahomes in the 2017 NFL Draft, which fundamentally elevated the Chiefs from a playoff-caliber team to a perennial Super Bowl contender.888 From 2013 to 2017, under coach Andy Reid but without starting quarterback Patrick Mahomes, the Chiefs made the playoffs in four of five seasons but won just one playoff game. From 2018 to 2023, the Chiefs had four Super Bowl appearances and won three. Another is Josh Allen, selected in 2018, who elevated the Buffalo Bills from years of ineptitude to one of the league’s top teams.999 From 2000 to 2017, the Bills made the playoffs once (in 2017). From 2018 to 2024, the Bills made the playoffs six times and made the AFC champsionship game twice. Under a different, but perhaps entirely reasonable, utility function––maximizing the probability of drafting transformational or elite players––decisions that appear suboptimal under Massey and Thaler’s framework may actually be entirely rational.
@@
A key insight, overlooked by previous researchers, is how variance in performance value changes over the course of the draft. Just as expected performance value declines convexly throughout the draft (see Figure 1a), so does variance (see Figure 3 in Section 4). We hypothesize that general managers are not only trading for expected surplus value but also for variance, which is directly related to the probability of acquiring a transformational player in the far right tail of performance. For example, consider the 2013 draft trade between the Dolphins and the Raiders. In this exchange, the Dolphins traded expected surplus value to the Raiders in exchange for variance. Based on the expected number of elite players,101010 Here, we define an elite player as one at least as valuable as Jared Goff, represented by the orange curve in Figure 14 in Section 6. the Dolphins’ draft capital increased by 40%. The Raiders traded a single pick worth 0.893 expected elite players for two picks collectively worth 0.639 (0.538 and 0.101). Although the Raiders accumulated substantially more expected surplus value, the Dolphins increased their expected number of elite players. Further, notice how much less the $42^{nd}$ pick is worth than before––it produces 89% fewer expected elite players than the $3^{rd}$ pick, but just 13% less expected surplus value. The likelihood of acquiring transformational players decays precipitously across the draft.
@@ 2 Measuring player performance
Since the value of a draft pick is realized by the on-field performance of the player it produces, a crucial first step is to measure the value of a player’s performance. Doing so takes care and consideration. Each play in American football consists of a complex interplay between 22 players of varying roles interacting in continuous time and space. This makes it difficult to measure how a player’s individual performance impacts the ultimate outcome of a play. For instance, a linebacker may impact a play simply by occupying space. He could cause the ball carrier to move in a different direction, and there is no easy way to measure the impact of that action. There is no industry-standard way of measuring player performance on a given play in American football, particularly in a way that measures players of all positions on the same scale. Hence, analysts use performance proxies. Each measure is imperfect and has its own drawbacks. We describe many of these performance measures in detail in Appendix B. Throughout this paper, for simplicity, we use one performance outcome variable, detailed below.
@@
Researchers, beginning with Massey and Thaler, (2013), have been interested in particular in the value of a player’s performance during his first contract. This is because an NFL player is paid particularly cheaply during his first contract. A player’s first contract is determined by his draft position and can be much lower than the value of his performance, creating a first contract “surplus value.” Teams can take advantage of the resulting extra salary cap space by signing good veteran players in free agency. The drawback of focusing solely on a player’s first contract is that the very best players stay with the same team across multiple contracts, providing value across their entire careers. For example, Tom Brady played in the NFL for 23 seasons and won seven Super Bowls. Continuing with the precedent set by Massey and Thaler, (2013) and subsequent analyses, in this study we consider just first contract performance and surplus value.
@@ 3 Traditional draft position value curves
Now, we review the traditional approach to constructing draft position value curves. Prior to the draft, we do not know how well a player drafted at draft position $x\in\{1,...,256\}$ will perform in the NFL. Thus, we think of the performance outcome $Y$ associated with position $x$ as a random variable, denoted by a capital letter. Recall that, throughout this paper, we let the realization of $Y$ be a player’s annualized second contract value as a percentage of the salary cap. Then, from the dataset of all recent draft picks in NFL history, we want to learn the de-noised relationship between pick number $x$ and performance outcome $Y$. Traditional analyses use a standard approach but with different measures of performance, detailed in Appendix B (Massey and Thaler,, 2013; Stuart,, 2012; Fitzgerald and Spielberger,, 2024; Pro Football Focus,, 2024; Baldwin,, 2024). The de-noised relationship they estimate from data is the expected value of $Y$ given $x$.
@@
Expected performance value is less steep than cost, indicating that players are underpaid on average in their first contracts. This produces positive expected surplus value across most of the draft (the long-dashed black line is mostly positive). Expected surplus value is larger than one throughout the first round, peaking in the middle of the first round. This surplus spike is the “loser’s curse” from Massey and Thaler, (2013): top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams.
The trade market curve (solid black line) is much steeper than the expected performance value curve (dotted black line) and the expected surplus value curve (long-dashed black line). Teams aren’t valuing draft positions by expected performance value or expected surplus value. The implication is that, if one adopts either of these utility functions, general managers are trading suboptimally: they don’t trade down often enough, and they often overpay when they trade up. According to expected surplus value, even teams at the very top of the draft should trade down, particularly towards the middle of the first round. This result is counterintuitive because very top picks are so highly regarded by general managers and football fans.

### v4

@@ Abstract
A longstanding question in the judgment and decision making literature is whether experts, even in high-stakes environments, exhibit the same cognitive biases observed in controlled experiments with inexperienced participants. Massey and Thaler (2013) claim to have found an example of bias and irrationality in expert decision making: general managers’ behavior in the National Football League draft pick trade market. They argue that general managers systematically overvalue top draft picks, which generate less surplus value on average than later first-round picks, a phenomenon known as the loser’s curse. Their conclusion hinges on the assumption that general managers should use expected surplus value as their utility function for evaluating draft picks. This assumption, however, is neither explicitly justified nor necessarily aligned with the strategic complexities of constructing a National Football League roster. In this paper, we challenge their framework by considering alternative utility functions, particularly those that emphasize the acquisition of transformational players––those capable of dramatically increasing a team’s chances of winning the Super Bowl. Under a decision rule that prioritizes the probability of acquiring elite players, which we construct from a novel Bayesian hierarchical Beta regression model, general managers’ draft trade behavior appears rational rather than systematically flawed. More broadly, our findings highlight the critical role of carefully specifying a utility function when evaluating the quality of decisions.
@@ 1 Introduction
A longstanding question in the judgment and decision making literature is whether experts, when making high-stakes decisions, exhibit the same cognitive biases and irrational behaviors documented in studies involving inexperienced participants in low-stakes environments. Critics of this literature argue that biases may be overstated, as most studies do not analyze decision making in real-world, high-stakes settings. Massey and Thaler, (2013) claim to have found an exception, applying behavioral economic principles to a context––the National Football League (NFL) draft––where decision makers, namely NFL general managers, are highly experienced, financially incentivized, and have access to extensive data and learning opportunities.
Massey and Thaler’s study has had a significant impact not just in sports analytics but in the broader decision making literature, reinforcing the notion that systematic biases can persist even in expert settings. They argue that general managers systematically overpay for top draft picks despite evidence that these selections generate less “surplus value”––the difference between the value of a player’s performance and his cost––than later first-round picks, on average. This so-called “loser’s curse,” widely cited as evidence of irrational decision making, is the subject of their ultimate striking assertion: “top draft picks are significantly overvalued in a manner that is inconsistent with rational expectations and efficient markets, and consistent with psychological research” (Massey and Thaler,, 2013).
@@ 1.1 Overview
Players entering the NFL are allocated to 32 teams through an annual draft. The draft consists of seven rounds, in which teams take turns selecting players in reverse order of the previous season’s win-loss records. The team with the worst record picks first, while the reigning Super Bowl champion selects last (with some minor exceptions). There are up to $256$ total draft picks because the NFL adds up to $32$ additional “compensatory free agent” draft picks––creating a hidden eigth round––that allow teams to replace players who they lost in free agency.111 https://operations.nfl.com/journey-to-the-nfl/the-nfl-draft/the-rules-of-the-draft Trading draft picks, involving the exchange of current and future draft selections as well as active players, is a fundamental component of constructing an NFL roster. Teams typically engage in these trades to secure specific players or improve positional depth. For example, in 2013 the Miami Dolphins traded the $3^{rd}$ overall pick to the Oakland Raiders in exchange for the $12^{th}$ and $42^{nd}$ overall picks.222 https://bleacherreport.com/articles/1618345-miami-dolphins-select-dion-jordan-with-no-3-pick-in-trade-with-oakland-raiders Trades like this raise fundamental questions about the valuation of draft picks. How did the Dolphins and Raiders determine that their trade was mutually beneficial? More broadly, if a team seeks to acquire a specific draft pick, what constitutes an appropriate offer? If presented with a trade proposal involving multiple draft picks, how should a team evaluate its value and decide whether to accept or reject the exchange? A rigorous analysis of draft pick valuation is essential to understanding the trade dynamics that shape NFL roster management.
@@
More recently, Massey and Thaler, (2013) developed less ad hoc draft position value curves, visualized in Figure 1. The solid line in Figure 1a denotes expected performance value, which quantifies the expected on-field performance of players drafted at a given position, expressed on a salary-equivalent scale. The dot-dashed line reflects first contract compensation, which denotes the cost of drafted players during their first contract. A player’s first contract compensation is essentially deterministic, as it is dictated by draft position. [^4] Drafted rookies may negotiate certain contract elements, such as performance bonuses or injury guarantees. However, this compensation is often much lower than the value of the player’s on-field performance, leading to what is known as first contract surplus value. The dashed lines in Figures 1a and 1b reflect expected surplus value, which measures the expected difference between a player’s first contract performance and his cost. Finally, the solid line in Figure 1b is the trade market curve, a draft position value curve that best fits the dataset of all observed trades involving only draft picks. This curve reflects the draft value implied by actual trade market behavior.
@@
Massey and Thaler asserted two main takeaways from their analysis displayed in Figure 1. First, they claimed to have found a loser’s curse, visualized in Figure 1a. Teams with the worst win-loss records in the previous season—the losers—receive the top draft picks, which produce higher performance on average since expected performance value declines monotonically and convexly throughout the draft. However, these teams are cursed because top draft picks generate less surplus value on average than mid-to-late first-round picks, which typically belong to better teams. Expected surplus value (i.e., expected performance above cost) increases throughout the early first round, peaking in the middle of the round. Second, Massey and Thaler asserted that, if one adopts the expected surplus value utility function, the draft pick trade market is highly inefficient, visualized in Figure 1b. The trade market curve is much steeper than the expected surplus value curve and differs in shape. Thus, teams consistently overpay to trade up in the draft, investing excessive draft capital to select “top draft picks [that] are significantly overvalued” (Massey and Thaler,, 2013). In contrast, trading down and accumulating additional draft picks is an effective strategy to maximize surplus value. For example, consider the 2013 draft trade between the Dolphins and the Raiders. Using expected surplus value,555 This calculation is based on our expected surplus value curve, visualized in Figure 2 in Section 3, as the exact values from Massey and Thaler’s curve are unavailable. the Raiders’ draft capital increased by 135%. The Dolphins traded two picks collectively worth 2.40 points (1.51 and 0.89, respectively) in exchange for a single pick worth 1.02. Notably, the $12^{th}$ overall pick had a higher expected surplus value than the $3^{rd}$ overall pick (i.e., $1.51>1.02$). Although the Dolphins won the trade according to Jimmy Johnson’s draft value chart, the Raiders accumulated substantially more expected surplus value.
@@
Massey and Thaler’s assertions, summarized in Figure 1, replicate to recent data (see Figure 2 in Section 3). Despite the widespread recognition of their work over an extended period, the draft strategies employed by general managers in the NFL have remained largely unchanged. This persistence raises an important question: why has there been no apparent adjustment in decision making? One possible explanation is that NFL general managers are resistant to change and continue to engage in suboptimal trade practices due to cognitive biases or misjudgment. However, this perspective prompts further inquiry: are these executives, responsible for managing multi-billion-dollar franchises, fundamentally misguided, or is there a crucial element missing from existing analyses? This paper explores the latter hypothesis, proposing an alternative explanation. Specifically, we argue that there may be a sound economic rationale underlying observed trade behavior. It is plausible that general managers are employing a different, yet entirely rational, utility function that does not rely solely on the expected future surplus value of a player at a given draft position.
@@
Perhaps the overarching strategic objective of NFL teams is to win the Super Bowl––a goal consistently emphasized by team owners, general managers, coaches, and players. As Kansas City Chiefs quarterback Patrick Mahomes succinctly stated, “the only goal is to win the Super Bowl.” [^6] https://x.com/_MLFootball/status/1826003252460269907?t=JvStTWyM92jEe21LxqOTPw&s=19 Massey and Thaler’s conclusions are predicated on the assumption that general managers, in pursuit of this objective, should adopt expected surplus value as their guiding utility function. However, while their analysis relies on this assumption, they do not explicitly acknowledge the primacy of winning the Super Bowl or establish a direct link between maximizing expected surplus value and achieving that goal. While it is possible that maximizing expected surplus value contributes to winning, alternative utility functions may yield entirely different strategic conclusions. This paper develops that possibility, exploring the idea that Super Bowl success hinges on acquiring transformational players––those capable of profoundly altering a team’s trajectory––rather than optimizing expected surplus value. One example is the selection of Patrick Mahomes in the 2017 NFL Draft, which fundamentally elevated the Chiefs from a playoff-caliber team to a perennial Super Bowl contender.777 From 2013 to 2017, under coach Andy Reid but without starting quarterback Patrick Mahomes, the Chiefs made the playoffs in four of five seasons but won just one playoff game. From 2018 to 2023, the Chiefs had four Super Bowl appearances and won three. Another is Josh Allen, selected in 2018, who elevated the Buffalo Bills from years of ineptitude to one of the league’s top teams.888 From 2000 to 2017, the Bills made the playoffs once (in 2017). From 2018 to 2024, the Bills made the playoffs six times and made the AFC champsionship game twice. Under a different, but perhaps entirely reasonable, utility function––maximizing the probability of drafting transformational or elite players––decisions that appear suboptimal under Massey and Thaler’s framework may actually be entirely rational.
@@
A key insight, overlooked by previous researchers, is how variance in performance value changes over the course of the draft. Just as expected performance value declines convexly throughout the draft (see Figure 1a), so does variance (see Figure 3 in Section 4). We hypothesize that general managers are not only trading for expected surplus value but also for variance, which is directly related to the probability of acquiring a transformational player in the far right tail of performance. For example, consider the 2013 draft trade between the Dolphins and the Raiders. In this exchange, the Dolphins traded expected surplus value to the Raiders in exchange for variance. Based on the expected number of elite players,999 Here, we define an elite player as one at least as valuable as Jared Goff, represented by the orange curve in Figure 14 in Section 6. the Dolphins’ draft capital increased by 40%. The Raiders traded a single pick worth 0.893 expected elite players for two picks collectively worth 0.639 (0.538 and 0.101). Although the Raiders accumulated substantially more expected surplus value, the Dolphins increased their expected number of elite players. Further, notice how much less the $42^{nd}$ pick is worth than before––it produces 89% fewer expected elite players than the $3^{rd}$ pick, but just 13% less expected surplus value. The likelihood of acquiring transformational players decays precipitously across the draft.
@@ 2 Measuring player performance
Since the value of a draft pick is realized by the on-field performance of the player it yields, a crucial first step is to measure the value of a player’s performance. Doing so takes care and consideration. Each play in American football consists of a complex interplay between 22 players of varying roles interacting in continuous time and space. This makes it difficult to measure how a player’s individual performance impacts the ultimate outcome of a play. For instance, a linebacker may impact a play simply by occupying space. He could cause the ball carrier to move in a different direction, and there is no easy way to measure the impact of that action. There is no industry-standard way of measuring player performance on a given play in American football, particularly in a way that measures players of all positions on the same scale. Hence, analysts use performance proxies. Each measure is imperfect and has its own drawbacks. We describe many of these performance measures in detail in Appendix B. Throughout this paper, for simplicity, we use one performance outcome variable, detailed below.
@@
Researchers, beginning with Massey and Thaler, (2013), have been interested in particular in the value of a player’s performance during his first contract. This is because an NFL player is paid particularly cheaply during his first contract. A player’s first contract is determined by his draft position and can be much lower than the value of his performance, creating a first contract “surplus value.” Teams can take advantage of the resulting extra salary cap space by signing good veteran players in free agency. The drawback of focusing solely on a player’s first contract is that the very best players stay with the same team across multiple contracts, providing value across their entire careers. For example, Tom Brady played in the NFL for 23 seasons and won seven Super Bowls, playing almost the entirety of his career with the team that drafted him. Continuing with the precedent set by Massey and Thaler, (2013) and subsequent analyses, in this study we consider just first contract performance and surplus value.
@@ 3 Traditional draft position value curves
We review the traditional approach to constructing draft position value curves. Prior to the draft, we do not know how well a player drafted at draft position $x\in\{1,...,256\}$ will perform in the NFL. Thus, we think of the performance outcome $Y$ associated with position $x$ as a random variable, denoted by a capital letter. Recall that, throughout this paper, we let the realization of $Y$ be a player’s annualized second contract value as a percentage of the salary cap. Then, from the dataset of all recent draft picks in NFL history, we want to estimate a de-noised relationship between pick number $x$ and performance outcome $Y$. Traditional analyses use a standard approach but with different measures of performance, detailed in Appendix B (Massey and Thaler,, 2013; Stuart,, 2012; Fitzgerald and Spielberger,, 2024; Pro Football Focus,, 2024; Baldwin,, 2024). The de-noised relationship they estimate from data is the expected value of $Y$ given $x$.
@@
Expected performance value is less steep than cost, indicating that players are underpaid on average in their first contracts. This produces positive expected surplus value across most of the draft (the long-dashed black line is mostly positive). Expected surplus value is larger than one throughout the first round, peaking in the middle of the first round. This surplus spike is what Massey and Thaler, (2013) assert is the “loser’s curse”: top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams.
The trade market curve (solid black line) is much steeper than the expected performance value curve (dotted black line) and the expected surplus value curve (long-dashed black line). This implies that teams aren’t valuing draft positions by expected performance value or expected surplus value. Alternatively, if general managers are in fact adopting either of these as the utility function, they are trading suboptimally: they don’t trade down often enough, and they often overpay when they trade up. According to expected surplus value, even teams at the very top of the draft should trade down, particularly towards the middle of the first round. This result is counterintuitive because very top picks are so highly regarded by general managers and football fans.